# Hitting probability for Reflected Brownian Motion at Small Target

### v1

@@ 2.1. Construction of RBM on General (SULD) Domains
thus $(Y^{x}_{t},L^{x}_{t})$ satisfies (1). As Brownian motion spend almost zero time on any point set we see that $\Delta=\{t:Y^{x}_{1}(t)=0\}$ is a Lebesgue null set. So by Proposition 2.2 we see that $(Y^{x}_{t},L^{x}_{t})$ is equal to $(X_{t}^{x},\xi_{t}^{x})$ with probability 1. ∎
@@
Suppose $\mathscr{V}=\left\{V^{0},V^{1},...\right\}$ is another cover of $\Omega$ defined in 2.4 and $(\Psi_{k})_{k\geq 1}$ be another collection of associated maps. Suppose for some $k$ and $m$ that $V_{m}\cap U_{k}$ is non-empty. Then the transition map
@@
2.2. Construction of RBM on Analytic Simply Connected Domains in $\mathbb{C}$
@@
Suppose $\Omega$ and Ωsuperscriptnormal-$\Omega^{\prime}$ are two simply connected open subset of $\mathbb{H}$ and $B^{z_{0}}_{t}$ is the RBM in $\mathbb{H}$ where $z_{0}\in\Omega$ with the associated local time process $\xi^{z_{0}}_{t}$. Let $f$ be a conformal automorphism of $\mathbb{H}$. Let $\tau_{\Omega}=\inf\{t\geq 0:B^{z_{0}}_{t}\in\partial\Omega\}$. Then with probability one, $(f(B^{z_{0}}_{t}),t\in[0,\tau))$ is equal to the time-changed RBM in $\mathbb{H}$.
@@
Let $f$ be a conformal map between $\Omega$ and $\mathbb{H}$. If $f$ can be extended conformally between a neighborhood Ωsuperscriptnormal-$\Omega^{\prime}$ of $\Omega$ and $\mathbb{H}^{\prime}$ of $\mathbb{H}$ with finite number of singularities on $\partial\Omega$, then $f$ maps the complex RBM in $\Omega$ to a time-changed complex RBM in $\mathbb{H}$.
@@ 5.1. Explicit Computation in $\mathbb{H}$
Define $h_{0}(z)=|\Im{z}|,h_{n}(z)=1-|1-|h_{n-1}(z)|$, and we notice that $h(z)=\lim_{n\to\infty}h_{n}(z)$ exists for all $z$. By proposition 5.3, $h$ works for the bijection. ∎
@@ 7. Appendix: Proofs of technical details
Suppose $\mathscr{V}=\left\{V^{0},V^{1},...\right\}$ is another cover of $\Omega$ defined in 2.4 and $(\Psi_{k})_{k\geq 1}$ be another collection of associated maps. Suppose for some $k$ and $m$ that $V_{m}\cap U_{k}$ is non-empty. Then the transition map
@@
Let $f$ be a conformal bijection between $\mathbb{D}$ and $\Omega$ where $\Omega$ is defined in 2.7. Then we may extend $f$ to a conformal bijection $F$ between Ωsuperscriptnormal-$\Omega^{\prime}$, an open neighborhood of $\Omega$ and $\mathbb{D}^{\prime}$, an open neighborhood of $\mathbb{D}$, and $F$ agree with $f$ on $\mathbb{D}$. [13] (Theorem 2.25)
@@
Suppose $\Omega$ and Ωsuperscriptnormal-$\Omega^{\prime}$ are two simply connected open subset of $\mathbb{H}$ and $B^{z_{0}}_{t}$ is the RBM in $\mathbb{H}$ where $z_{0}\in\Omega$ with the associated local time process $\xi^{z_{0}}_{t}$. Let $f$ be a conformal automorphism of $\mathbb{H}$. Let $\tau_{\Omega}=\inf\{t\geq 0:B^{z_{0}}_{t}\in\partial\Omega\}$. Then with probability one, $(f(B^{z_{0}}_{t}),t\in[0,\tau))$ is equal to the time-changed RBM in $\mathbb{H}$.

### v2

@@ 2.1. Construction of RBM on General (SULD) Domains
where $Z^{x}_{1}$ is a Brownian motion. Thus $(Y^{x}_{t},L^{x}_{t})$ satisfies (1). As Brownian motion spend almost zero time on any point set we see that $\Delta=\{t:Y^{x}_{1}(t)=0\}$ is a Lebesgue null set. So by Proposition 2.2 we see that $(Y^{x}_{t},L^{x}_{t})$ is equal to $(X_{t}^{x},\xi_{t}^{x})$ with probability 1. ∎
@@
Remark 2.4.
By proposition 2.3 we see the reflected Brownian motion in $\mathbb{H}_{n}$ is a semi-martingale.
@@
Suppose $\mathscr{V}=\left\{V^{0},V^{1},...\right\}$ is another cover of $\Omega$ defined in 2.5 and $(\Psi_{k})_{k\geq 1}$ be another collection of associated maps. Suppose for some $k$ and $m$ that $V_{m}\cap U_{k}$ is non-empty. Then the transition map
@@
2.2. Equivalent Construction of RBM on Analytic Simply Connected Domains in $\mathbb{C}$
@@
Suppose $\Omega$ and $\Omega^{\prime}$ are two simply connected open subset of $\mathbb{H}$ and $B^{z_{0}}_{t}$ is the RBM in $\mathbb{H}$ where $z_{0}\in\Omega$ with the associated local time process $\xi^{z_{0}}_{t}$. Let $f$ be a conformal automorphism of $\mathbb{H}$. Let $\tau_{\Omega}=\inf\{t\geq 0:B^{z_{0}}_{t}\in\partial\Omega\}$. Then with probability one, $(f(B^{z_{0}}_{t}),t\in[0,\tau))$ is equal to the time-changed RBM in $\mathbb{H}$.
@@
Let $f$ be a conformal map between $\Omega$ and $\mathbb{H}$. If $f$ can be extended conformally between a neighborhood $\Omega^{\prime}$ of $\Omega$ and $\mathbb{H}^{\prime}$ of $\mathbb{H}$ with finite number of singularities on $\partial\Omega$, then $f$ maps the complex RBM in $\Omega$ to a time-changed complex RBM in $\mathbb{H}$.
@@ 5.1. Explicit Computation in $\mathbb{H}$
Define $h_{0}(z)=|\Im{z}|,h_{n}(z)=1-|1-|h_{n-1}(z)||$, and we notice that $h(z)=\lim_{n\to\infty}h_{n}(z)$ exists for all $z$. By proposition 5.3, $h$ works for the bijection. ∎
@@ 7. Appendix: Proofs of technical details
Suppose $\mathscr{V}=\left\{V^{0},V^{1},...\right\}$ is another cover of $\Omega$ defined in 2.5 and $(\Psi_{k})_{k\geq 1}$ be another collection of associated maps. Suppose for some $k$ and $m$ that $V_{m}\cap U_{k}$ is non-empty. Then the transition map
@@
Let $f$ be a conformal bijection between $\mathbb{D}$ and $\Omega$ where $\Omega$ is defined in 2.8. Then we may extend $f$ to a conformal bijection $F$ between $\Omega^{\prime}$, an open neighborhood of $\Omega$ and $\mathbb{D}^{\prime}$, an open neighborhood of $\mathbb{D}$, and $F$ agree with $f$ on $\mathbb{D}$. [13] (Theorem 2.25)
@@
Suppose $\Omega$ and $\Omega^{\prime}$ are two simply connected open subset of $\mathbb{H}$ and $B^{z_{0}}_{t}$ is the RBM in $\mathbb{H}$ where $z_{0}\in\Omega$ with the associated local time process $\xi^{z_{0}}_{t}$. Let $f$ be a conformal automorphism of $\mathbb{H}$. Let $\tau_{\Omega}=\inf\{t\geq 0:B^{z_{0}}_{t}\in\partial\Omega\}$. Then with probability one, $(f(B^{z_{0}}_{t}),t\in[0,\tau))$ is equal to the time-changed RBM in $\mathbb{H}$.